OpenStudy (twistrose):

Please tell me how to solve this and tell me the answer on how to simplify this ratio: 16 to 56?

OpenStudy (anonymous):

16 : 56 is same as 16/56, you simplify it as you woud do with fraction

OpenStudy (anonymous):

let's say you have 2:4, that is same as 2/4 = 1/2 so the simplified ratio would be 1:2

OpenStudy (twistrose):

Can you show me how? I've tried it before but I can't figure it out.

OpenStudy (twistrose):

Thank you

OpenStudy (anonymous):

just write here your answer so i can tell you if it's correct :)

rebeccaxhawaii (rebeccaxhawaii):

both are divisible by 8 so 16/8=2 56/8=7 so 2:7 is your final answer

OpenStudy (twistrose):

I got 4 to 14, and that was by dividing it by four, but since I had to divide by two again I thought that I had to divide by 6, so I did but that is wrong.

OpenStudy (anonymous):

okay and 4/14 can still be divided with 2 and that is your answer :) btw, you did it great by yourself :D

OpenStudy (twistrose):

oh thanks :)

OpenStudy (anonymous):

you're welcome